# Generalized finite and affine $W$-algebras in type $A$

### v1

@@ 3.1 Affine vertex algebra and free fermion vertex algebra
Let $e$ be the nilpotent element in $\mathfrak{g}$ associated with the pyramid $\lambda$ as defined in (2.2). As before, we denote by $\mathfrak{a}$ be the centralizer of $e$ in $\mathfrak{g}.$ We will use the invariant bilinear form (|)(\ |\ )( | ) on $\mathfrak{a}$ which was introduced in [2]. Specifically, set $\text{deg}_{\lambda}(e_{ij})=\text{col}_{\lambda}(j)-\text{col}_{\lambda}(i)$ to induce the $\mathbb{Z}$-gradations
@@
To use a version of the BRST complex of the quantum Drinfeld–Sokolov reduction, associated with the Lie algebra $\mathfrak{a}$ by analogy with [8, Ch. 15], and extending the construction of [14], introduce two vertex algebras as follows. The first is $V^{k}(\mathfrak{a})$, the affine vertex algebra at the level $k\in\mathbb{C}$. It is defined as
@@
$[a\otimes t^{n},b\otimes t^{m}]=[a,b]\otimes t^{n+m}+\delta_{n+m,0}(a|b)K$ (3.2)
@@
$[E_{i_{1},j_{1}}^{(r_{1})}(z),E_{i_{2},j_{2}}^{(r_{2})}(w)]=[E_{i_{1},j_{1}}^{(r_{1})},E_{i_{2},j_{2}}^{(r_{2})}](w)\delta(z,w)+k(E_{i_{1},j_{1}}^{(r_{1})}\ |\ E_{i_{2},j_{2}}^{(r_{2})})\partial_{w}\delta(z,w),$
@@
$[\phi_{\mathsf{n}}{}_{\lambda}\phi^{\mathsf{m}}]=[\phi^{\mathsf{m}}{}_{\lambda}\phi_{\mathsf{n}}]={\mathsf{m}}({\mathsf{n}})$
@@ 3.2 Definition of the generalized affine $W$-algebras
    Observe that $[d{}_{\lambda}a]=\sum_{I\in S_{\lambda,\mu}}[\,:\phi^{I}E_{I}:\,{}_{\lambda}\,a\,].$ By skew-symmetry, we have
@@
    $\sum_{I\in S_{\lambda,\mu}}\big{[}:\phi^{I}E_{I}:{}_{\lambda}\ \phi_{\mathsf{n}}\big{]}$ $=\sum_{I\in S_{\lambda,\mu}}:[\phi_{\mathsf{n}\ -\lambda-\partial}\ \phi^{I}]\,E_{I}:=\mathsf{n}.$
@@
    $:\displaystyle=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:[\phi_{\mathsf{n}\ -\lambda-\partial}\phi^{I}]:\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}::-\sum_{I,I^{\prime}\,\in S_{\lambda,\mu}}:\phi^{I}[\phi_{\mathsf{n}\ -\lambda-\partial}:\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:]$ :
@@
    $[d_{\lambda}\phi_{\mathsf{n}}]=\mathsf{n}+\chi(\mathsf{n})+\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[E_{I},\mathsf{n}]}:\ .$
@@
    $[d{}_{\lambda}\phi^{\mathsf{m}}]$ $::\displaystyle=\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}\big{[}\phi^{\mathsf{m}}_{\ \lambda}\phi_{[E_{I^{\prime}},E_{I}]}\big{]}:$
@@
To verify the relation $Q^{2}=0$, it is enough to show that $[d{}_{\lambda}d]=0$. We have
@@
$::\displaystyle=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\frac{1}{2}::\phi^{I^{\prime}}\phi^{E_{I^{\prime}}\cdot E_{I}^{\ast}}:E_{I}:-\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]:$
$::\displaystyle=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\frac{1}{2}::\phi^{I}\phi^{I^{\prime}}:[E_{I^{\prime}},E_{I}]:-\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]:$
@@
One can easily check that $[d_{\lambda}\phi^{\chi}]=0$ and hence it remains to verify that
@@
$::\displaystyle=\frac{1}{4}\sum_{I,I^{\prime},I^{\prime\prime},I^{\prime\prime\prime}\in S_{\lambda,\mu}}(E_{I})^{\ast}\left([E_{I^{\prime\prime\prime}},E_{I^{\prime\prime}}]\right)::\phi^{I^{\prime\prime}}\phi^{I^{\prime\prime\prime}}:\ :\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:$ (3.7)
@@
    If $B\in\ker Q$ and $A\in C^{k}(\lambda,\mu)$ then $[Q(A)_{\lambda}B]$ is in $\mathbb{C}[\lambda]\otimes\text{im}(Q)$.
@@
Suppose $Q(B)=A$ for some $B\in C^{k}(\lambda,\mu).$ Since $Q(\partial B)=\partial A,$ the first assertion follows. The second and third assertions follow from (3.14) and (3.15), respectively, by taking into account the relation $[A_{\lambda}Q(B)]=0.$
@@ 3.3 Structure of the generalized affine $W$-algebra
$[J_{a}{}_{\lambda}\phi_{\mathsf{n}}]=\phi_{[a,\mathsf{n}]},\quad[J_{a}{}_{\lambda}\phi^{\mathsf{m}}]=\phi^{a\cdot\mathsf{m}}\ ,$ (3.18)
@@
$[J_{a}{}_{\lambda}J_{b}]=J_{[a,b]}+(k(a|b)+\text{tr}_{\mathfrak{n}_{\lambda,\mu}}(\pi_{+}\text{ad }a\circ\pi_{+}\text{ad }b))\lambda.$ (3.20)
@@
$H(C^{k}(\lambda,\mu),Q)\simeq H(C_{+}^{k}(\lambda,\mu),Q|_{C_{+}^{k}(\lambda,\mu)})\otimes H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})\simeq H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q}),$
@@
The generalized $W$-algebra $W^{k}(\lambda,\mu)$ is isomorphic to $H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q}).$
@@
$\text{gr}(\phi^{I})$ $=\left(\text{deg}(E_{I})+\frac{1}{2},-\text{deg}(E_{I})+\frac{1}{2}\right).$
@@
The next lemma concerns the properties of the complex $(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})$. To state it, recall that $r_{-}=\phi^{\mathfrak{n}_{\lambda,\mu}^{*}}\oplus J_{\mathfrak{p}}$ and set
@@
In the complex $(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})$ we have the following.
@@
Invoking [6, Lemma 4.18 and Theorem 4.19], we come to the following theorem.
@@
$H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})=H^{0}(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})\simeq V(H(R_{-},\widetilde{Q}))$
where $V(H(R_{-},\widetilde{Q}))$ is the universal enveloping vertex algebra of $H(R_{-},\widetilde{Q}).$
@@
$W^{k}(\lambda,\mu)\ \ {\text{\em is a vertex subalgebra of }}\ \ V(J_{\mathfrak{p}}),$ (3.28)
where $V(J_{\mathfrak{p}})$ is the universal enveloping vertex algebra of $\mathbb{C}[\partial]\otimes J_{\mathfrak{p}}$ endowed with the $\lambda$-bracket introduced in (3.20). Take an ordered basis $\mathcal{B}_{\mathfrak{p}}$ of $J_{\mathfrak{p}}$. Then $V(J_{\mathfrak{p}})$ has the PBW basis induced from the ordered set ~=$\widetilde{\mathcal{B}}=\{\,\partial^{t}\mathcal{B}_{\mathfrak{p}}\,|\,t\in\mathbb{Z}_{+}\,\}$ and we can define degrees of elements of $V(J_{\mathfrak{p}})$ by degrees with respect to the basis elements in $\widetilde{\mathcal{B}}$. In particular, the linear term of any element will mean the degree $1$ part in $V(J_{\mathfrak{p}})$ with respect to the given basis of $J_{\mathfrak{p}}$.
@@
By Theorem 3.10, the number of free generators of $W^{k}(\lambda,\mu)$ is dim(ker $\text{dim}(\text{ker }\varphi)$. Lemma 3.8 implies that
dim(ker $\text{dim}(\text{ker }\varphi)=\text{dim }\mathfrak{p}-\text{dim }\mathfrak{n}_{\lambda,\mu}^{\ast}=\text{dim }\mathfrak{p}-\text{dim }\mathfrak{n}_{\lambda,\mu}=\text{dim }\mathfrak{a}(0).$
@@
In the particular case $\lambda=(1^{N})$, our BRST complex defining the $W$-algebra $W^{k}(\lambda,\mu)$ coincides with the one used in [6, Sec. 5] to describe the $W$-algebra $W^{k}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{gl}_{N}$ of type $\mu$. Therefore, these two affine $W$-algebras are the same. ∎
@@ 4.1 Zhu algebra of $W^{k}(\lambda,\mu)$
$[Zhu_{H}(a),Zhu_{H}(b)]_{\hbar}=\sum_{j\in\mathbb{Z}_{+}}{\Delta_{a}-1\choose j}\,\hbar^{j}\,Zhu_{H}(a_{(j)}b).$
@@
Note that $[h,a]=j_{a}a$ for $a\in\mathfrak{a}(j_{a})$. In other words, the grading $\bigoplus_{m\in\mathbb{Z}}\mathfrak{a}(m)$ can be induced from the Dynkin grading by $h$ on $\mathfrak{gl}_{N}$. Using (4.2), we obtain
$[\bar{a},\bar{b}]=[Zhu_{H}(a),Zhu_{H}(b)]=Zhu_{H}([a,b])+(\Delta_{a}-1)(a|b)k=Zhu_{H}([a,b])-j_{a}(a|b)k$
@@
$[\bar{a},\bar{b}]=\overline{[a,b]},\quad[\bar{\phi}^{\mathsf{m}},\bar{\phi}_{\mathsf{n}}]=\mathsf{m}(\mathsf{n}).$
@@
On the other hand, since the Hamiltonian operator $H$ on $C^{k}(\lambda,\mu)$ induces the Hamiltonian operator on the affine $W$-algebra $W^{k}(\lambda,\mu)$, we get the associative algebra $Zhu_{H}(W^{k}(\lambda,\mu)).$ By applying [6, Theorem 4.20], we come to the following proposition.
@@
Furthermore, due to Lemma 3.9, the assumptions of [6, Theorem 4.20] are satisfied. Thus, there is a canonical associative algebra isomorphism

### v2

@@ 3.1 Affine vertex algebra and free fermion vertex algebra
Let $e$ be the nilpotent element in $\mathfrak{g}$ associated with the pyramid $\lambda$ as defined in (2.2). As before, we denote by $\mathfrak{a}$ be the centralizer of $e$ in $\mathfrak{g}.$ We will use the invariant bilinear form (|)(\ |\ ) on $\mathfrak{a}$ which was introduced in [4]. Specifically, set $\text{deg}_{\lambda}(e_{ij})=\text{col}_{\lambda}(j)-\text{col}_{\lambda}(i)$ to induce the $\mathbb{Z}$-gradations
@@
Remark 3.1.
We can consider another bilinear form induced from that of $\mathfrak{gl}_{N}=\mathfrak{sl}_{N}\oplus\mathbb{C}.$ The main results in this paper such as Theorem 3.12, Corollary 3.13, and Theorem 4.6 are still true even when this bilinear form is used. However, we follow the convention used in [4], [25], and [24].
To use a version of the BRST complex of the quantum Drinfeld–Sokolov reduction, associated with the Lie algebra $\mathfrak{a}$ by analogy with [15, Ch. 15], and extending the construction of [24], introduce two vertex algebras as follows. The first is $V^{k}(\mathfrak{a})$, the affine vertex algebra at the level $k\in\mathbb{C}$. It is defined as
@@
$a\otimes t^{n},b\otimes t^{m}]=[a,b]\otimes t^{n+m}+\delta_{n+m,0}(a|b)K$ (3.2)
@@
$E_{i_{1},j_{1}}^{(r_{1})}(z),E_{i_{2},j_{2}}^{(r_{2})}(w)]=[E_{i_{1},j_{1}}^{(r_{1})},E_{i_{2},j_{2}}^{(r_{2})}](w)\delta(z,w)+k(E_{i_{1},j_{1}}^{(r_{1})}\ |\ E_{i_{2},j_{2}}^{(r_{2})})\partial_{w}\delta(z,w),$
@@
$\phi_{\mathsf{n}}{}_{\lambda}\phi^{\mathsf{m}}]=[\phi^{\mathsf{m}}{}_{\lambda}\phi_{\mathsf{n}}]={\mathsf{m}}({\mathsf{n}})$
@@ 3.2 Definition of the generalized affine $W$-algebras
    Observe that $d{}_{\lambda}a]=\sum_{I\in S_{\lambda,\mu}}[\,:\phi^{I}E_{I}:\,{}_{\lambda}\,a\,].$ By skew-symmetry, we have
@@
    $\sum_{I\in S_{\lambda,\mu}}\big[:\phi^{I}E_{I}:{}_{\lambda}\ \phi_{\mathsf{n}}\big]$ $=\sum_{I\in S_{\lambda,\mu}}:[\phi_{\mathsf{n}\ -\lambda-\partial}\ \phi^{I}]\,E_{I}=\mathsf{n}.$
@@
    $=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:[\phi_{\mathsf{n}\ -\lambda-\partial}\phi^{I}]:\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}::-\sum_{I,I^{\prime}\,\in S_{\lambda,\mu}}:\phi^{I}[\phi_{\mathsf{n}\ -\lambda-\partial}:\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:]:$
@@
    $d_{\lambda}\phi_{\mathsf{n}}]=\mathsf{n}+\chi(\mathsf{n})+\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[E_{I},\mathsf{n}]}:\ .$
@@
    $[d{}_{\lambda}\phi^{\mathsf{m}}]$ $=\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}\big[\phi^{\mathsf{m}}_{\ \lambda}\phi_{[E_{I^{\prime}},E_{I}]}\big]::$
@@
To verify the relation $Q^{2}=0$, it is enough to show that $d{}_{\lambda}d]=0$. We have
@@
$=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\frac{1}{2}::\phi^{I^{\prime}}\phi^{E_{I^{\prime}}\cdot E_{I}^{\ast}}:E_{I}:-\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]::$
$=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\frac{1}{2}::\phi^{I}\phi^{I^{\prime}}:[E_{I^{\prime}},E_{I}]:-\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]::$
@@
One can easily check that $d_{\lambda}\phi^{\chi}]=0$ and hence it remains to verify that
@@
$=\frac{1}{4}\sum_{I,I^{\prime},I^{\prime\prime},I^{\prime\prime\prime}\in S_{\lambda,\mu}}(E_{I})^{\ast}\left([E_{I^{\prime\prime\prime}},E_{I^{\prime\prime}}]\right)::\phi^{I^{\prime\prime}}\phi^{I^{\prime\prime\prime}}:\ :\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}::$ (3.7)
@@
    If $B\in\ker Q$ and $A\in C^{k}(\lambda,\mu)$ then $Q(A)_{\lambda}B]$ is in $\mathbb{C}[\lambda]\otimes\text{im}(Q)$.
@@
Suppose $Q(B)=A$ for some $B\in C^{k}(\lambda,\mu).$ Since $Q(\partial B)=\partial A,$ the first assertion follows. The second and third assertions follow from (3.14) and (3.15), respectively, by taking into account the relation $A_{\lambda}Q(B)]=0.$
@@ 3.3 Structure of the generalized affine $W$-algebra
$J_{a}{}_{\lambda}\phi_{\mathsf{n}}]=\phi_{[a,\mathsf{n}]},\quad[J_{a}{}_{\lambda}\phi^{\mathsf{m}}]=\phi^{a\cdot\mathsf{m}}\ ,$ (3.18)
@@
$J_{a}{}_{\lambda}J_{b}]=J_{[a,b]}+(k(a|b)+\text{tr}_{\mathfrak{n}_{\lambda,\mu}}(\pi_{+}\text{ad }a\circ\pi_{+}\text{ad }b))\lambda.$ (3.20)
@@
[…]
@@
[…]
@@
$\text{gr}(\phi^{I})$ $=\left(-\text{deg}(E_{I})+\frac{1}{2},\text{deg}(E_{I})+\frac{1}{2}\right).$
@@
[…]
@@
[…]
@@
Remark 3.11.
The choice of (2.10) was made so as to ensure Lemma 3.9 and Lemma 3.10 (3). Let $\chi:\mathfrak{n}_{\lambda,\mu}\to\mathbb{C}$ be an arbitrary character. The condition $\chi(E_{i,i+1}^{(\lambda_{i+1}-1)})\neq 0$ is necessary for Lemma 3.9. If one chooses $\chi$ different from (2.10) but still satisfying this condition, then Lemma 3.10 (3) and Theorem 3.12 remain valid.
Invoking [9, Lemma 4.18 and Theorem 4.19], we come to the following theorem.
@@
$)H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})=H^{0}(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})\simeq V(H(R_{-},\widetilde{Q}$))
[…]
@@
$W^{k}(\lambda,\mu)\ \ {\text{is a vertex subalgebra of }}\ \ V(J_{\mathfrak{p}}),$ (3.28)
where $V(J_{\mathfrak{p}})$ is the universal enveloping vertex algebra of $\mathbb{C}[\partial]\otimes J_{\mathfrak{p}}$ endowed with the $\lambda$-bracket introduced in (3.20). Take an ordered basis $\mathcal{B}_{\mathfrak{p}}$ of $J_{\mathfrak{p}}$. Then $V(J_{\mathfrak{p}})$ has the PBW basis induced from the ordered set $\widetilde{\mathcal{B}}=\{\,\partial^{t}\mathcal{B}_{\mathfrak{p}}\,|\,t\in\mathbb{Z}_{+}\,\}$ and we can define degrees of elements of $V(J_{\mathfrak{p}})$ by degrees with respect to the basis elements in $\widetilde{\mathcal{B}}$. In particular, the linear term of any element will mean the degree $1$ part in $V(J_{\mathfrak{p}})$ with respect to the given basis of $J_{\mathfrak{p}}$.
@@
By Theorem 3.12, the number of free generators of $W^{k}(\lambda,\mu)$ is $\text{dim}(\text{ker }\varphi)$. Lemma 3.9 implies that
$\text{dim}(\text{ker }\varphi)=\text{dim }\mathfrak{p}-\text{dim }\mathfrak{n}_{\lambda,\mu}^{\ast}=\text{dim }\mathfrak{p}-\text{dim }\mathfrak{n}_{\lambda,\mu}=\text{dim }\mathfrak{a}(0).$
@@
  1. (a)
    In the particular case $\lambda=(1^{N})$, our BRST complex defining the $W$-algebra $W^{k}(\lambda,\mu)$ coincides with the one used in [9, Sec. 5] to describe the $W$-algebra $W^{k}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{gl}_{N}$ of type $\mu$. Therefore, these two affine $W$-algebras are the same.
  2. (b)
    For other types of Lie algebras, similar construction should work. The only nontrivial part is finding various gradations on $\mathfrak{a}=\mathfrak{g}^{e},$ which substitute the partition $\mu$ in $\mathfrak{g}=\mathfrak{gl}_{N}$ case.
@@ 4.1 Zhu algebra of $W^{k}(\lambda,\mu)$
$Zhu_{H}(a),Zhu_{H}(b)]_{\hbar}=\sum_{j\in\mathbb{Z}_{+}}{\Delta_{a}-1\choose j}\,\hbar^{j}\,Zhu_{H}(a_{(j)}b).$
@@
Note that $h,a]=j_{a}a$ for $a\in\mathfrak{a}(j_{a})$. In other words, the grading $\bigoplus_{m\in\mathbb{Z}}\mathfrak{a}(m)$ can be induced from the Dynkin grading by $h$ on $\mathfrak{gl}_{N}$. Using (4.2), we obtain
$\bar{a},\bar{b}]=[Zhu_{H}(a),Zhu_{H}(b)]=Zhu_{H}([a,b])+(\Delta_{a}-1)(a|b)k=Zhu_{H}([a,b])-j_{a}(a|b)k$
@@
$\bar{a},\bar{b}]=\overline{[a,b]},\quad[\bar{\phi}^{\mathsf{m}},\bar{\phi}_{\mathsf{n}}]=\mathsf{m}(\mathsf{n}).$
@@
On the other hand, since the Hamiltonian operator $H$ on $C^{k}(\lambda,\mu)$ induces the Hamiltonian operator on the affine $W$-algebra $W^{k}(\lambda,\mu)$, we get the associative algebra $Zhu_{H}(W^{k}(\lambda,\mu)).$ By applying [9, Theorem 4.20], we come to the following proposition.
@@
Furthermore, due to Lemma 3.10, the assumptions of [9, Theorem 4.20] are satisfied. Thus, there is a canonical associative algebra isomorphism